Boundary Condition Implementation

• Implicit BCs are used in PUMA when any of the implicit time-integration schemes are used (i.e. the BC is evaluated at the next timestep as opposed to the current timestep).

• The boundary condition Jacobian J_{BC} (\frac{∂Q_{BC}}{∂Q}) is obtained numerically using central differences, which is completely general and does much to improve the maintainability of the code.

• There are several types of BCs implemented in PUMA, and more can be added without much difficulty as and when wanted.

Type Description  Variable values specified
0  Do nothing
1  Fixed at freestream
2 Fixed at given values ρ,u,v,w,p
3  First order extrapolation from the interior
4  Second order extrapolation from the interior
5  Inflow/outflow (Riemann)
6  Specified p_{0}, T_{0}  p_{0}, T_{0}
7  Fixed back pressure pressure  p
8  Tangency
9  No slip adiabatic wall
10 No slip fixedtemperature wall T_{wall}  T_{wall}
11 Zero u velocity
12 Zero v velocity
13 Zero w velocity